The perimiter of a rectangular field is 228 yards. If the width of the field is 52 yards, what is the length?
Answer by TimothyLamb(4379)

2L + 2W = 228
W = 52
---
2L + 2W = 228
2L + 2(52) = 228
2L + 104 = 228
2L = 124
---
L = 62
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php